How to solve:
If triangular area is 160ft^2 and the height is 4ft more than it's base. Find the base.

Answer by richwmiller(17219)

1/2 b*h=a
b*h=320
h=b+4,
(b+4)*b=320
b^2+4b -320=0
b = 16, h = 20